OpenStudy (richyw):

improper integral. sorry for all of these. my textbook thinks that "converges" is a suitable explanation. \[\int^{\infty}_0 \tan\frac{1}{x}\text{d}x\]

OpenStudy (richyw):

I need to say if it converges or diverges.

OpenStudy (abb0t):

Take the integral as you would from [0, R]

OpenStudy (anonymous):

can you change variables to \(u=\frac{1}{x}\) and then flip the limits of integration?

OpenStudy (abb0t):

Then take the limit of the function as R->∞

OpenStudy (richyw):

is this sufficient to say it converges? \[\lim_{x \to \infty} \tan\frac{1}{x}=0\]

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

thing thing has an infinite number of singularities

OpenStudy (anonymous):

how about the substitution \[x = \tan^{-1}u\]

OpenStudy (richyw):

oh right.

OpenStudy (anonymous):

or \[x=\cot^{-1}(u)\]

OpenStudy (anonymous):

the problem isn't just that the path is infinite,

OpenStudy (anonymous):

it is undefined infinitely often

OpenStudy (klimenkov):

What about \(\tan\frac1x\sim\frac1x,\, x\rightarrow\infty\) ?

OpenStudy (anonymous):

yep!! convergence....

OpenStudy (anonymous):

i think there is a mistake here there is no way this thing converges

OpenStudy (anonymous):

\[\lim_{x\to0}\tan(1/x)=\infty\]

OpenStudy (richyw):

oh wait a second, it does diverge.

OpenStudy (anonymous):

graph it and see what it looks like, then tell me it converges...

OpenStudy (anonymous):

@satellite73 can we say that as \(x\to0\), it is infinite or is it undefined?

OpenStudy (richyw):

but the limits iof integration are supposed to be from 1 to infty. so ignore the zero

OpenStudy (anonymous):

makes no difference, the problem is not the infinite path tangent is undefined infinitely often

OpenStudy (klimenkov):

Use that \(\tan\frac1x\sim\frac1x,\, x\rightarrow\infty\) and immediately say that it diverges.

OpenStudy (anonymous):

@richyw it also has infinite "vertical asymptotes" and hence it is discontinuous over the interval and cannot be integrated!

OpenStudy (anonymous):

@satellite73 so, you are saying that the function cannot exist to begin with?

OpenStudy (richyw):

@klimenkov is that just the first term of the taylor expansion at infinity?

OpenStudy (klimenkov):

If you want Taylor series - yes. But from \(\tan x\sim x,\,x\rightarrow0\) implies \(\tan\frac1x\sim\frac1x,\, x\rightarrow\infty\).

OpenStudy (richyw):

how am I supposed to know that \(\tan x \sim x,\;x\to 0\) ?

OpenStudy (klimenkov):

Compute\[\lim_{x\rightarrow0}\frac{\tan x}{x}\]

OpenStudy (richyw):

AWSOME. ok this is what I was struggling with last night.

OpenStudy (richyw):

why does doing that show that they behave the same?

OpenStudy (richyw):

that step is what really confuses me!

OpenStudy (richyw):

same step was used in the second line of this. I don't understand why I do this!

OpenStudy (klimenkov):

Try to check this theorem in your calculus book.

OpenStudy (richyw):

believe me I have tried. I even emailed my prof asking where the theorem is. does it have a name?

OpenStudy (klimenkov):

Ok. Did you understand why \(\tan\frac1x\sim\frac1x,\, x\rightarrow\infty\) ?

OpenStudy (richyw):

no, because I don't understand why \(\tan x \sim x,\;x\to 0\)

OpenStudy (klimenkov):

Do you know that \[\lim_{x\rightarrow0}\frac{\sin x}{x}=1\]

OpenStudy (richyw):

yes

OpenStudy (klimenkov):

Now \[\lim_{x\rightarrow0}\frac{\tan x}{x}=\lim_{x\rightarrow0}\frac{\frac{\sin x}{\cos x}}{x}=\lim_{x\rightarrow0}\frac{\sin x}{x}=1\]because \(\lim_{x\rightarrow0}\cos x=1\). Got it?

OpenStudy (richyw):

yes, I can compute the limit.

OpenStudy (richyw):

but I don't understand its significance

OpenStudy (klimenkov):

Ah. Ok. If \(f(x)\sim g(x),\,x\rightarrow a\), it means that \(f(x)\) behaves like \(g(x)\) at the point \(x=a\).

OpenStudy (richyw):

right. do you know why?

OpenStudy (richyw):

I mean why the limit thing, not that definition

OpenStudy (klimenkov):

Or \[\lim_{x\rightarrow a}\frac{f(x)}{g(x)}=1\]

OpenStudy (richyw):

AHHH that's so obvious!

OpenStudy (klimenkov):

It is equivalent definitions.

OpenStudy (richyw):

omg. thank you so much.

OpenStudy (richyw):

I feel so stupid now, but if you looked at my questions history I spent like 3 hours last night trying to get this explanation!

OpenStudy (richyw):

thank you so much!

OpenStudy (klimenkov):

Don't be shy to ask. It is normal not to see obvious things because obvious is a relative concept. I wish you a lot of luck in learning math. Love it.